OpenStudy (wesdg1978):

How do you find the value of df^-1/dx at x = f(a)? f(x) = 5x + 10, a = 4

OpenStudy (anonymous):

f(x) = 5x + 10 f'(x) = 5 at x = f(4), f(4) = 5(4) + 10

OpenStudy (wesdg1978):

That's what I thought too, but the answer choices are 1/5 or 5 or 10 or 1/10

OpenStudy (anonymous):

f'(x) = 5